The hypotenuse of a right triangle measures 7√2 inches. One of the legs measures 2√7 inches. Find the measure of the other leg.

Answer by mananth(16946)

hyp^2 = leg1^2 +leg2^2
(7sqrt2)^2= (2sqrt7)^2 + leg2^2
49*2=4*7+leg2^2
91-28= leg2^2
63= leg2^2
3*3*7 = leg2^2
3sqrt7= leg2
Ananth